OpenStudy (lukecrayonz):

Find the area under the graph of f over the interval [5,8] 8x+3 for x<6 57-x for x>6

OpenStudy (lukecrayonz):

it is 8x+3 for x \[8x+3, x \le 6\]

OpenStudy (anonymous):

The integral of a function represents the area under it. \[Area = \int\limits_{5}^{6}(8x+3)dx + \int\limits_{6}^{8}(57-x)dx\]

OpenStudy (lukecrayonz):

So then you integrate and solve for a number, and I keep ending with an expression.

OpenStudy (anonymous):

i don't understand

OpenStudy (lukecrayonz):

A statement :P It should be a number! Idk what im doing wrong

OpenStudy (anonymous):

to solve the first part it is F(6) - F(5) second part [after plus sign] F(8) - F(6)

OpenStudy (anonymous):

\[\int\limits_{5}^{6}(8x+3)dx = 4x^2 + 3x\]\[\int\limits_{6}^{8}(57 - x) dx = 57x - \frac{ x^2 }{ 2 }\]

OpenStudy (anonymous):

first part: \[4(6)^2 + 3(6) - (4(5)^2 + 3(5))\] seond part:\[57(8) - \frac{ 8^2 }{ 2 } - (57(6) - \frac{6^2}{2})\]

OpenStudy (anonymous):

i get 47 + 100 = 147

OpenStudy (lukecrayonz):

It's correct! I have no idea what I was doing wrong lol, I think it was just a brain fart.

OpenStudy (lukecrayonz):

Thank you so much, highly appreciate it. Nobody likes to answer my questions anymore haha

OpenStudy (anonymous):

glad i could help ^_^ might be because the smartscores & medals have been not working for a few days now. i get it demotivated people :P

OpenStudy (anonymous):

i guess*

OpenStudy (lukecrayonz):

@Euler271 could you quickly do this?\[\int\limits_{0}^{4}(5e^t)dt\]

OpenStudy (lukecrayonz):

Im getting 14899 and it's marking wrong

OpenStudy (anonymous):

\[5\int\limits_{0}^{4}e^t = 5e^t \left(\begin{matrix}4 \\ 0\end{matrix}\right)= 5 (e^4 - e^0) = 5(e^4 - 1)\]

OpenStudy (anonymous):

e^t is the derivative/integral of itself 5 is just a constant so we don't consider it in integration

OpenStudy (lukecrayonz):

Right! that's what I got, so why am I getting such a weird number? What is the value of 5(e^4-1)?

OpenStudy (anonymous):

267.9907502 but i doubt they want an approximation. you should literally type 5(e^4 - 1) in [if this is online homework]

OpenStudy (lukecrayonz):

They got 6699.77 lol :(

OpenStudy (lukecrayonz):

Lol sometimes my online book is wrong..

OpenStudy (lukecrayonz):

Area under the graph of f over the interval [-3,3] 2-x^2 if x<0 2 if x>(greater than or equal to) 0

OpenStudy (anonymous):

lol yup. textbooks have mistakes sometimes. i've driven myself crazy because of those.

OpenStudy (anonymous):

\[Area = \int\limits_{-3}^{0} (2 - x^2)dx + \int\limits_{0}^{3}2dx\]

OpenStudy (lukecrayonz):

I GOT -9 AND IT'S NOT AN ANSWERRRRRRRRRRRRRR

OpenStudy (anonymous):

\[2x - \frac{ x^3 }{ 3 }\left(\begin{matrix}0 \\ -3\end{matrix}\right) + 2x \left(\begin{matrix}3 \\ 0\end{matrix}\right)\]\[= 2(0) - \frac{ 0^3 }{ 3 } - (2(-3) - \frac{ (-3)^3 }{ 3 }) + 2(3) - 2(0)\] area can't be negative ^_^ i'll calculate this 2 sec

OpenStudy (lukecrayonz):

LOL not an answer still ;_; 9 3 -15 -6

OpenStudy (lukecrayonz):

OH so its 9

OpenStudy (lukecrayonz):

:O???

OpenStudy (anonymous):

it's 3. my bad

OpenStudy (anonymous):

so many negative signs

OpenStudy (lukecrayonz):

Okay! I'll trust your work over mine since I'm new to this. Find the area of the shaded region f(x)=-x^3-x^2+20x, g(x)=0 [-5,4]

OpenStudy (anonymous):

can you make a rough sketch of the shaded region plz?

OpenStudy (anonymous):

since some of it is below the x-axis, doing it normally would lead to whatever is below the x-axis being negative (i.e subtracted). so you must consider them seperately and add them so you need to do: \[\int\limits_{-5}^{0}(-x^3-x^2+20x) dx + \int\limits_{0}^{4}(-x^3-x^2+20x) dx\] the answer: http://www.wolframalpha.com/input/?i=integrate+-x%5E3-x%5E2%2B20x+from+-5+to+4 you should definitely try it though ^_^

OpenStudy (anonymous):

sorry, wolfram is giving you the integral (which can be negative) however, the AREA itself must be positive, so: absolute value of this: (positive): http://www.wolframalpha.com/input/?i=integrate+-x%5E3-x%5E2%2B20x+from+-5+to+0 plus this: http://www.wolframalpha.com/input/?i=integrate+-x%5E3-x%5E2%2B20x+from+0+to+4

OpenStudy (lukecrayonz):

That's not an option D: lol

OpenStudy (lukecrayonz):

-81/12 2137/12 1625/12 -679/12

OpenStudy (anonymous):

weird.